OpenStudy (anonymous):

Josephine started a business selling cosmetics. She spent $4500 to obtain her merchandise, and it costs her $200 per week for general expenses. She earns $550 per week in sales. What is the minimum number of weeks it will take for Josephine to make a profit? A. 13 weeks B. 14 weeks C. 12 weeks D. 15 weeks

OpenStudy (anonymous):

4,500 + 200 (per week)= 4,700 550 (per week) - 200 (per week)= 350 (per week) 4,700 (divided by) 350= 14 weeks @CrazyCountryGirl